# Learning Structured Representations by Embedding Class Hierarchy with Fast Optimal Transport

### v1

@@ 4 Experiments
For a fair comparison, we closely aligned our implementation with Zeng et al. (2022) to ensure consistency. This section will cover dataset benchmarks, the multi-modality of features, downstream hierarchical classification retrieval, and interpretability analysis. A detailed description of training hyperparameters is provided in App.E, along with additional experiments, including the effect of label hierarchy structure, label hierarchy tree weights, optimal transport flow settings (distribution of $a$ in Eq.3), and comparisons with non-CPCC hierarchical methods, in App.G. Throughout this section, the Flat objective is the standard cross entropy loss in Eq.1 with $\lambda=0$.
@@ 4.1 Datasets
We conducted experiments across $7$ diverse datasets, CIFAR10, CIFAR100 (Krizhevsky and Hinton, 2009), BREEDS benchmarks (Santurkar et al., 2021) including four settings LIVING17 (L17), ENTITY13 (E13), ENTITY30 (E30), NONLIVING26 (N26), and iNaturalist-mini (Van Horn et al., 2018) (INAT) to thoroughly evaluate the efficacy and influence of OT-CPCC methods. The dataset statistics is shown in Tab.3. Note that BREEDS and INAT are larger scale datasets with high resolution images. As subsets of ImageNet, BREEDS inherit a more complicated label hierarchy and more realistic, possibly multi-modal class distributions. The major difference from ImageNet is that BREEDS’s class hierarchy is manually calibrated so that it reflects the visual hierarchy instead of the WordNet semantic hierarchy in the original ImageNet, which removes ill-defined hierarchical relationships unrelated to vision tasks.
@@ 4.2 Existence of Multi-mode Class-conditioned Feature Distribution
To verify our hypothesis in Fig.1, we aim to determine whether class-conditioned clusters exhibit a multi-modal structure in practice. We track the class-conditioned features during training: at each epoch, for each class-conditioned feature cluster, we fit a Gaussian Mixture Model (GMM) with the number of components ranging from 1 to 10. We report the optimal number of components based on the lowest Akaike Information Criterion (AIC), averaged across all fine-level class labels. If the best number of Gaussian components exceeds 1, this may indicate the presence of a multi-modal structure in the feature distributions, assuming a sufficiently large batch size. From Fig.6, with default training hyperparameters in App.E, we observe that this phenomenon occurs across most datasets for both $\ell_{2}$-CPCC and OT-CPCC training, suggesting that using optimal transport distances in CPCC computation is more appropriate.
@@ 4.3 Hierarchical Classification and Retrieval
Table 4: Generalization performance on coarse level. The prefix of metric represents the split of test set: s = source, t = target. The average number of three seeds are reported. See the detailed version including standard deviation in Tab.14,15 and performance on each BREEDS setting is in Tab.16,17. Due to the computation complexity, we do not report EMD results on training BREEDS from scratch.
@@
For downstream classification and retrieval, we split each dataset into source and target subsets, with both further divided into train and test sets, as illustrated in Fig.4. The source and target datasets share the same level of coarse labels but differ in fine labels, where the fine level refers to the more granular class labels that are exactly one level below the coarse labels. For example, in CIFAR10, the source set includes the fine labels (deer, dog, frog, horse) and (ship, truck), while the target set includes (bird, cat) and (airplane, automobile). Both sets share the same coarse labels: animal and transportation. We include details of hierarchical construction and splits for other datasets in App.D.
@@ 4.4 Interpretability Evaluation
A model is more interpretable when its representations accurately capture hierarchical information. CPCC is a natural measure of the precision of the hierarchical structure embedded in corresponding metric spaces. Unlike rank-based metrics, where only the relative order of distances matters, CPCC not only requires the correct ranking of distances (i.e., fine classes sharing the same coarse class should be closer than fine classes with different coarse parents), but is also influenced by the absolute values of the node-to-node distances, because CPCC is affected by the given label tree weights (see App.G.2 for more discussion). Therefore, we report CPCC, a stricter metric, on the test set. In Tab.3, we present $\ell_{2}$-CPCC for the Flat and $\ell_{2}$ objectives, and use the corresponding OT distances for the OT-CPCC experiments.
@@ 5 Related Work
We discuss methods that operate on feature representations, thereby directly influencing the geometry of the representation space when hierarchical information is provided. DeViSE (Frome et al., 2013) is an early multimodal work to align image and text embedding evaluated on a hierarchical precision metric to reflect semantic knowledge in the representation. Ge et al. (2018); Zhang et al. (2016) modify triplet loss to capture class similarity for metric learning or image retrieval. Barz and Denzler (2019); Garg et al. (2022) embed classes on hypersphere based on hierarchical similarity, and similar ideas have been explored in the hyperbolic space (Kim et al., 2023) where tree can be embedded with low distortion. Yang et al. (2022) provides a multi-task proxy loss for deep metric learning. Nolasco and Stowell (2022) proposes a ranking loss to make Euclidean distance of feature close to a target $d$-dim sequence determined by the rank of tree metric. We work on the extension of Zeng et al. (2022) which is a special type of hierarchical methods where a regularizer is used. For a broader discussion on various hierarchical methods, we direct readers to the summary provided by Bertinetto et al. (2020).
@@ 6 Conclusion
Acknowledgement
SZ and HZ are partially supported by an NSF IIS grant No. 2416897. HZ would like to thank the support from a Google Research Scholar Award. MY was supported by MEXT KAKENHI Grant Number 24K03004. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies.
@@ Appendix A Approximation Methods of Earth Mover’s Distance
Most approximation methods do not rely on any tree structure. Sinkhorn [Cuturi, 2013] is a fully differentiable approximation method which leverages the entropic regularization setup of EMD and the convergence of doubly stochastic matrix. Sliced Wasserstein Distance (SWD) [Bonneel et al., 2015] projects data to 1 dimension through a random projection matrix and uses $\text{EMD}_{\text{1d}}$ for fast calculation, but the quality of approximation highly relies on the number of random projections. Tree Wasserstein Distance (TWD) [Yamada et al., 2022, Le et al., 2019, Takezawa et al., 2021] is a tree based approximation algorithm that can be written in the closed matrix form $\left\lVert\operatorname{diag}(\mathrm{w})\bm{B}(a-b)\right\rVert_{1}$, where $\mathrm{w}$ contains weights of ${\mathcal{T}}$ and ${\bm{B}}$ is a 0-1 internal nodes by leaf nodes matrix that represents if a node is the ancestor of a leaf node. To make sure TWD is a good approximation, an extra step of lasso regression is performed to learn $\mathrm{w}$ such that the tree metric approximates Euclidean distance between features with iterative gradient based method. In Yamada et al. [2022], it requires us to learn the tree distance between all pairs of nodes, which leads to quadratic computational complexity with respect to sample size as shown in Tab.1.
Appendix B EMD-CPCC Generalizes $\ell_{2}$-CPCC
@@
Dowson and Landau [1982] showed the following analytical form of ${\mathcal{W}}_{2}^{2}$: $\lVert\mu_{z}-\mu_{z^{\prime}}\rVert^{2}+\text{tr}\left(\Sigma_{z}+\Sigma_{z^{\prime}}-2(\Sigma_{z}\Sigma_{z^{\prime}})^{1/2}\right)$. The trace term becomes 00 if $\Sigma_{z}=\Sigma_{z^{\prime}}=\Sigma$.
Appendix C Details for Fast Flow Tree CPCC Algorithm
In this section, we first present the full pseudocode to compute FastFT-CPCC in Alg.3. Compared to FlowTree [Backurs et al., 2020] which uses the helper function in Alg.2 [Kalantari and Kalantari, 1995], we replace the greedy_flow_matching algorithm with Alg.1 which aligns with algorithm that computes the 1 dimensional OT transportation plan.
@@ Appendix C Details for Fast Flow Tree CPCC Algorithm
Alg.1, or the flow matrix solution for 1-$d$ EMD problem, can be used when Monge Property or north-west-corner rules [Hoffman, 1963] hold. It can be shown by induction that this $O(m+n)$ greedy algorithm can be applied whenever the following constraint is satisfied: if $\{(z_{i},a_{i})\},\{(z_{j},b_{j})\}$ are indexed (1d feature, weight) sequences, for indices $i_{1}<i_{2}$ and $j_{1}<j_{2}$:
@@
Appendix D Dataset Hierarchy Construction
In this section, we discuss the details of dataset hierarchies and how we created the dataset splits. Throughout the main body of this paper, we consider a two-level hierarchy: coarse and fine. For each coarse node, except for BREEDS, we split the coarse node’s fine children into approximately 40% for the target set and 60% for the source set. Following this procedure, the splits are detailed in Tables 6, 7, 8, 9, 10, 11, and 12. The original class labels for CIFAR10 and CIFAR100 are from https://www.cs.toronto.edu/~kriz/cifar.html, for iNaturalist from https://github.com/visipedia/inat_comp/tree/master/2021, and for BREEDS from WordNet [Miller, 1995].
@@ Appendix D Dataset Hierarchy Construction
In Zeng et al. [2022], fine classes in BREEDS are defined one level deeper than the original labels. However, due to the large number of classes and limited samples per class within a batch, the distinction between $\ell_{2}$-CPCC and OT-CPCC can become less apparent. To address this, we employ a three-level hierarchy, using the labels from Santurkar et al. [2021] as fine labels and setting their parent categories as mid and coarse labels. Consequently, due to BREEDS setting, the fine classes are the same for both source and target datasets, so there is no source-target split in Tables 9, 10, 11, and 12. Despite this hierarchical structure, as seen in Fig.6, BREEDS exhibits relatively low multi-modality. With this design, in Sec.4.3, for classification and retrieval experiments, one-shot fine-tuning is not required for BREEDS on the target split.
@@
Appendix E Training Hyperparameters
@@ E.3 INAT
Appendix F Hierarchical Baseline Objectives
@@
We adopt all hierarchical baselines from the comparative study in Zeng et al. [2022], including:
Multi-Task Learning (MTL) [Caruana, 1997], where the backbone model is trained with two linear heads—one for coarse classes and one for fine classes—each with a cross-entropy loss. The total loss, $\mathcal{L}_{\textnormal{MTL}}$, is the sum of these two cross-entropy terms, with a scaling coefficient of 1 for both. Curriculum Learning (Curr) [Bengio et al., 2009] involves two training phases: first, training a coarse head for 50 epochs, followed by training the full network with a fine head for 200 epochs, starting from the coarse-trained features, assuming that coarse classification is an easier task. Sum Loss [Hu et al., 2018] uses only one fine classifier head, where coarse predictions are obtained by summing the fine class probabilities for the same coarse parent and applying a second cross-entropy loss. Hierarchical Cross Entropy (HXE) [Bertinetto et al., 2020] is the sum of weighted cross-entropy losses ($\alpha=0.4$), where the hierarchical structure is incorporated into the log-probabilities. The same paper also introduces Soft Label (Soft), which uses a single softmax loss based on the height of the least common ancestor between label pairs, with a hyperparameter $\beta=10$. Quadruplet Loss (Quad) [Zhang et al., 2016] extends triplet loss by defining two types of positive labels based on the hierarchy. For detailed mathematical expressions of these objectives, refer to App.C in Zeng et al. [2022].
@@
(SHC) [Garg et al., 2022] is the combination of MTL and Sum Loss. For a two level hierarchy, they jointly train a two-headed network. Define the network as $h:\mathcal{X}\rightarrow\Delta_{k_{1}}\times\Delta_{k_{2}}$ with a shared feature encoder $f:\mathcal{X}\rightarrow\mathcal{Z}$ and two classifiers $g_{\textnormal{coarse}}:\mathcal{Z}\rightarrow\Delta_{k_{1}},g_{\textnormal{fine}}:\mathcal{Z}\rightarrow\Delta_{k_{2}}$, with $k_{1}$ as the number of coarse classes and $k_{2}$ the number of fine classes. The ground truth coarse probability $p_{\text{coarse}}=g_{\textnormal{coarse}}(f(\cdot))$ is the output of coarse head after softmax normalization, and $\hat{p}_{\text{coarse}}$ is derived from the output of fine head. Let $\mathbf{W}$ be a $k_{1}$ by $k_{2}$ matrix representing the relationships in the label tree: if a fine class $i$ belongs to a coarse class $j$, then $\mathbf{W}_{ji}$ is $1$, else the entry is set to 00. Then $\hat{p}_{\text{coarse}}=\mathbf{W}g_{\text{fine}}(f(\cdot))$, so the probability of belonging to a given coarse class is the sum of the probabilities of its fine class descendants in the tree. Then, the SHC loss is defined as the Jensen-Shannon Divergence between these two probability distributions:
@@
In Garg et al. [2022], $\mathcal{L}_{\text{SHC}}$ is used in the combination with Flat cross entropy loss. Therefore, we treat it as a counterpart of CPCC losses and use $\mathcal{L}_{\textnormal{Flat}}+\lambda\cdot\mathcal{L}_{\textnormal{SHC}}$ as the training loss, with $\lambda=1$.
@@
(GC) [Garg et al., 2022] is also computed under the MTL two-head pipeline. The coarse classifier can be decomposed to a linear layer and softmax function, i.e., let $\mathbf{W}^{c}\in\mathbb{R}^{d\times k_{1}}$, $g_{\text{coarse}}(\cdot)=\textnormal{softmax}(\mathbf{W}^{c}(\cdot))$, and we define $\mathbf{W}^{f}$ similarly for the linear fine layer.
@@
(Rank) [Nolasco and Stowell, 2022] proposes a ranking loss to make Euclidean distance of feature $\rho$ close to the target $d$-dim sequence determined by the rank of tree metric:
@@
Appendix G Additional Experiments
G.1 Effect of Tree Depth

### v2

@@ 4 Experiments
For a fair comparison, we closely aligned our implementation with Zeng et al. (2022) to ensure consistency. This section will cover dataset benchmarks, the multi-modality of features, downstream hierarchical classification retrieval, and interpretability analysis. A detailed description of training hyperparameters is provided in App.F, along with additional experiments, including the effect of label hierarchy structure, label hierarchy tree weights, optimal transport flow settings (distribution of $a$ in Eq.3), and comparisons with non-CPCC hierarchical methods, in App.H. Throughout this section, the Flat objective is the standard cross entropy loss in Eq.1 with $\lambda=0$ without tree information.
@@ 4.1 Datasets
We conducted experiments across $7$ diverse datasets, CIFAR10, CIFAR100 (Krizhevsky & Hinton, 2009), BREEDS benchmarks (Santurkar et al., 2021) including four settings LIVING17 (L17), ENTITY13 (E13), ENTITY30 (E30), NONLIVING26 (N26), and iNaturalist-mini (Van Horn et al., 2018) (INAT) to thoroughly evaluate the efficacy and influence of OT-CPCC methods. The dataset statistics is shown in Tab.3. Note that BREEDS and INAT are larger scale datasets with high resolution images. As subsets of ImageNet, BREEDS inherit a more complicated label hierarchy and more realistic, possibly multi-modal class distributions. The major difference from ImageNet is that BREEDS’s class hierarchy is manually calibrated so that it reflects the visual hierarchy instead of the WordNet semantic hierarchy in the original ImageNet, which removes ill-defined hierarchical relationships unrelated to vision tasks.
@@ 4.2 Existence of Multi-mode Class-conditioned Feature Distribution
To verify our hypothesis in Fig.1, we aim to determine whether class-conditioned clusters exhibit a multi-modal structure in practice. We track the class-conditioned features during training: at each epoch, for each class-conditioned feature cluster, we fit a Gaussian Mixture Model (GMM) with the number of components ranging from 1 to 10. We report the optimal number of components based on the lowest Akaike Information Criterion (AIC), averaged across all fine-level class labels. If the best number of Gaussian components exceeds 1, this may indicate the presence of a multi-modal structure in the feature distributions, assuming a sufficiently large batch size. From Fig.6, with default training hyperparameters in App.F, we observe that this phenomenon occurs across most datasets for both $\ell_{2}$-CPCC and OT-CPCC training, suggesting that using optimal transport distances in CPCC computation is more appropriate.
@@ 4.3 Hierarchical Classification and Retrieval
Table 4: Generalization performance on coarse level. The prefix of metric represents the split of test set: s = source, t = target. The average number of three seeds are reported. See the detailed version including standard deviation in Tab.16,17 and performance on each BREEDS setting is in Tab.18,19. Due to the computation complexity, we do not report EMD results on training BREEDS from scratch.
@@
For downstream classification and retrieval, we split each dataset into source and target subsets, with both further divided into train and test sets, as illustrated in Fig.4. The source and target datasets share the same level of coarse labels but differ in fine labels, where the fine level refers to the more granular class labels that are exactly one level below the coarse labels. For example, in CIFAR10, the source set includes the fine labels (deer, dog, frog, horse) and (ship, truck), while the target set includes (bird, cat) and (airplane, automobile). Both sets share the same coarse labels: animal and transportation. We include details of hierarchical construction and splits for other datasets in App.E.
@@ 4.4 Interpretability Evaluation
A model is more interpretable when its representations accurately capture hierarchical information. CPCC is a natural measure of the precision of the hierarchical structure embedded in corresponding metric spaces. Unlike rank-based metrics, where only the relative order of distances matters, CPCC not only requires the correct ranking of distances (i.e., fine classes sharing the same coarse class should be closer than fine classes with different coarse parents), but is also influenced by the absolute values of the node-to-node distances, because CPCC is affected by the given label tree weights. Therefore, we report CPCC, a stricter metric, on the test set. In Tab.3, we present $\ell_{2}$-CPCC for the Flat and $\ell_{2}$ objectives, and use the corresponding OT distances for the OT-CPCC experiments.
@@ 5 Related Work
We discuss methods that operate on feature representations, thereby directly influencing the geometry of the representation space when hierarchical information is provided. DeViSE (Frome et al., 2013) is an early multimodal work to align image and text embedding evaluated on a hierarchical precision metric to reflect semantic knowledge in the representation. Ge et al. (2018); Zhang et al. (2016) modify triplet loss to capture class similarity for metric learning or image retrieval. Barz & Denzler (2019); Garg et al. (2022) embed classes on hypersphere based on hierarchical similarity, and similar ideas have been explored in the hyperbolic space (Kim et al., 2023) where tree can be embedded with low distortion. Yang et al. (2022) provides a multi-task proxy loss for deep metric learning. Nolasco & Stowell (2022) proposes a ranking loss to make Euclidean distance of feature close to a target $d$-dim sequence determined by the rank of tree metric. We work on the extension of Zeng et al. (2022) which is a special type of hierarchical methods where a regularizer is used. For a broader discussion on various hierarchical methods, we direct readers to the summary provided by Bertinetto et al. (2020).
@@ 6 Conclusion
Acknowledgment
SZ and HZ are partially supported by an NSF IIS grant No. 2416897. HZ would like to thank the support from a Google Research Scholar Award. MY is partly supported by MEXT KAKENHI Grant Number 24K03004 and by JST ASPIRE JPMJAP2302. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies.
@@ Appendix A Approximation Methods of Earth Mover’s Distance
Most approximation methods do not rely on any tree structure. Sinkhorn (Cuturi, 2013) is a fully differentiable approximation method which leverages the entropic regularization setup of EMD and the convergence of doubly stochastic matrix. Sliced Wasserstein Distance (SWD) (Bonneel et al., 2015) projects data to 1 dimension through a random projection matrix and uses $\text{EMD}_{\text{1d}}$ for fast calculation, but the quality of approximation highly relies on the number of random projections. Tree Wasserstein Distance (TWD) (Yamada et al., 2022; Le et al., 2019; Takezawa et al., 2021) is a tree based approximation algorithm that can be written in the closed matrix form $\left\lVert\operatorname{diag}(\mathrm{w})\bm{B}(a-b)\right\rVert_{1}$, where $\mathrm{w}$ contains weights of ${\mathcal{T}}$ and ${\bm{B}}$ is a 0-1 internal nodes by leaf nodes matrix that represents if a node is the ancestor of a leaf node. To make sure TWD is a good approximation, an extra step of lasso regression is performed to learn $\mathrm{w}$ such that the tree metric approximates Euclidean distance between features with iterative gradient based method. In Yamada et al. (2022), it requires us to learn the tree distance between all pairs of nodes, which leads to quadratic computational complexity with respect to sample size as shown in Tab.1.
Appendix B Comparison of Structured Representation Learning and Learning Label Representations
Generally speaking, the problem of structured representation learning and the problem of learning label representations, or learning label embeddings, are two different topics. In label representation learning, each label is mapped to a high-dimensional vector for downstream tasks. For example, in standard multi-class classification, we typically use one-hot vector embeddings. Alternative approaches include using a soft label to assign probability values in (0,1) to each class (Nguyen et al., 2014), or learning label embeddings (Weston et al., 2011; Sun et al., 2017; Liu et al., 2024). Label representation learning methods have various downstream applications, such as text summarization (Sun et al., 2017), novelty detection (Barz & Denzler, 2019), and extreme classification (Jalan & Kar, 2019; Tagami, 2017; Bhatia et al., 2015; Yu et al., 2014; Evron et al., 2018; Gupta et al., 2019).
While $\ell_{2}$-CPCC falls into this category, OT-CPCC methods are not label embedding methods. Instead of representing a class with a Euclidean centroid in $\ell_{2}$, OT-CPCC methods intentionally avoid this step by considering a more nuanced relationship between labels. OT-CPCC relies on pairwise relationships between data points from two classes to include more complex distributional geometry during learning. Since these methods depend on data point relations, there is no one-to-one correspondence between a label and a representation.
Appendix C EMD-CPCC Generalizes $\ell_{2}$-CPCC
@@
Dowson & Landau (1982) showed the following analytical form of
${\mathcal{W}}_{2}^{2}=\lVert\mu_{z}-\mu_{z^{\prime}}\rVert^{2}+\text{tr}\left(\Sigma_{z}+\Sigma_{z^{\prime}}-2(\Sigma_{z}\Sigma_{z^{\prime}})^{1/2}\right)$ (7)
This distance metric is also sometimes called Bures-Wasserstein (Bures) distance. The trace term becomes 00 if $\Sigma_{z}=\Sigma_{z^{\prime}}=\Sigma$.
Appendix D Details for Fast Flow Tree CPCC Algorithm
In this section, we first present the full pseudocode to compute FastFT-CPCC in Alg.3. Compared to FlowTree (Backurs et al., 2020) which uses the helper function in Alg.2 (Kalantari & Kalantari, 1995), we replace the greedy_flow_matching algorithm with Alg.1 which aligns with algorithm that computes the 1 dimensional OT transportation plan.
@@ Appendix D Details for Fast Flow Tree CPCC Algorithm
Proof Sketch
We can observe that the blue code blocks, which contain the update rule of the flow matrix, are almost identical between Alg.1 and Alg.2. According to Kalantari & Kalantari (1995), Alg.2 provides the optimal flow matrix for the optimal transport problem using the tree metric as the ground metric in ${\bm{D}}$. From Fig.1, the tree distance between any two samples in the augmented tree is always the same, meaning the order of encountering any vertex within the source or target distribution does not matter. This effectively reduces the FlowTree problem in Alg.2 to a flat 1d EMD problem in Alg.1. The major difference between two algorithms is that, unlike Alg.2, Alg.1 does not require passing the flow from the leaf to the root node.
@@
Alg.1, or the flow matrix solution for 1-$d$ EMD problem, can be used when Monge Property or north-west-corner rules (Hoffman, 1963) hold. It can be shown by induction that this $O(m+n)$ greedy algorithm can be applied whenever the following constraint is satisfied: if $\{(z_{i},a_{i})\},\{(z_{j},b_{j})\}$ are indexed (1d feature, weight) sequences, for indices $i_{1}<i_{2}$ and $j_{1}<j_{2}$:
@@
Appendix E Dataset Hierarchy Construction
In this section, we discuss the details of dataset hierarchies and how we created the dataset splits. Throughout the main body of this paper, we consider a two-level hierarchy: coarse and fine. For each coarse node, except for BREEDS, we split the coarse node’s fine children into approximately 40% for the target set and 60% for the source set. Following this procedure, the splits are detailed in Tables 6, 7, 8, 9, 10, 11, and 12. The original class labels for CIFAR10 and CIFAR100 are from https://www.cs.toronto.edu/~kriz/cifar.html, for iNaturalist from https://github.com/visipedia/inat_comp/tree/master/2021, and for BREEDS from WordNet (Miller, 1995).
@@ Appendix E Dataset Hierarchy Construction
In Zeng et al. (2022), fine classes in BREEDS are defined one level deeper than the original labels. However, due to the large number of classes and limited samples per class within a batch, the distinction between $\ell_{2}$-CPCC and OT-CPCC can become less apparent. To address this, we employ a three-level hierarchy, using the labels from Santurkar et al. (2021) as fine labels and setting their parent categories as mid and coarse labels. Consequently, due to BREEDS setting, the fine classes are the same for both source and target datasets, so there is no source-target split in Tables 9, 10, 11, and 12. Despite this hierarchical structure, as seen in Fig.6, BREEDS exhibits relatively low multi-modality. With this design, in Sec.4.3, for classification and retrieval experiments, one-shot fine-tuning is not required for BREEDS on the target split.
@@
Appendix F Training Hyperparameters
@@ F.3 INAT
Appendix G Hierarchical Baseline Objectives
@@
We adopt all hierarchical baselines from the comparative study in Zeng et al. (2022), including:
Multi-Task Learning (MTL) (Caruana, 1997), where the backbone model is trained with two linear heads—one for coarse classes and one for fine classes—each with a cross-entropy loss. The total loss, $\mathcal{L}_{\textnormal{MTL}}$, is the sum of these two cross-entropy terms, with a scaling coefficient of 1 for both. Curriculum Learning (Curr) (Bengio et al., 2009) involves two training phases: first, training a coarse head for 50 epochs, followed by training the full network with a fine head for 200 epochs, starting from the coarse-trained features, assuming that coarse classification is an easier task. Sum Loss (Hu et al., 2018) uses only one fine classifier head, where coarse predictions are obtained by summing the fine class probabilities for the same coarse parent and applying a second cross-entropy loss. Hierarchical Cross Entropy (HXE) (Bertinetto et al., 2020) is the sum of weighted cross-entropy losses ($\alpha=0.4$), where the hierarchical structure is incorporated into the log-probabilities. The same paper also introduces Soft Label (Soft), which uses a single softmax loss based on the height of the least common ancestor between label pairs, with a hyperparameter $\beta=10$. Quadruplet Loss (Quad) (Zhang et al., 2016) extends triplet loss by defining two types of positive labels based on the hierarchy. For detailed mathematical expressions of these objectives, refer to App.C in Zeng et al. (2022).
@@
(SHC) (Garg et al., 2022) is the combination of MTL and Sum Loss. For a two level hierarchy, they jointly train a two-headed network. Define the network as $h:\mathcal{X}\rightarrow\Delta_{k_{1}}\times\Delta_{k_{2}}$ with a shared feature encoder $f:\mathcal{X}\rightarrow\mathcal{Z}$ and two classifiers $g_{\textnormal{coarse}}:\mathcal{Z}\rightarrow\Delta_{k_{1}},g_{\textnormal{fine}}:\mathcal{Z}\rightarrow\Delta_{k_{2}}$, with $k_{1}$ as the number of coarse classes and $k_{2}$ the number of fine classes. The ground truth coarse probability $p_{\text{coarse}}=g_{\textnormal{coarse}}(f(\cdot))$ is the output of coarse head after softmax normalization, and $\hat{p}_{\text{coarse}}$ is derived from the output of fine head. Let $\mathbf{W}$ be a $k_{1}$ by $k_{2}$ matrix representing the relationships in the label tree: if a fine class $i$ belongs to a coarse class $j$, then $\mathbf{W}_{ji}$ is $1$, else the entry is set to 00. Then $\hat{p}_{\text{coarse}}=\mathbf{W}g_{\text{fine}}(f(\cdot))$, so the probability of belonging to a given coarse class is the sum of the probabilities of its fine class descendants in the tree. Then, the SHC loss is defined as the Jensen-Shannon Divergence between these two probability distributions:
@@
In Garg et al. (2022), $\mathcal{L}_{\text{SHC}}$ is used in the combination with Flat cross entropy loss. Therefore, we treat it as a counterpart of CPCC losses and use $\mathcal{L}_{\textnormal{Flat}}+\lambda\cdot\mathcal{L}_{\textnormal{SHC}}$ as the training loss, with $\lambda=1$.
@@
(GC) (Garg et al., 2022) is also computed under the MTL two-head pipeline. The coarse classifier can be decomposed to a linear layer and softmax function, i.e., let $\mathbf{W}^{c}\in\mathbb{R}^{d\times k_{1}}$, $g_{\text{coarse}}(\cdot)=\textnormal{softmax}(\mathbf{W}^{c}(\cdot))$, and we define $\mathbf{W}^{f}$ similarly for the linear fine layer.
@@
(Rank) (Nolasco & Stowell, 2022) proposes a ranking loss to make Euclidean distance of feature $\rho$ close to the target $d$-dim sequence determined by the rank of tree metric:
@@
Appendix H Additional Experiments
H.1 Effect of Tree Depth